OpenStudy (anonymous):

Find x - 3y if x = 3 and y = -2.

OpenStudy (anonymous):

ur answer will be 9

OpenStudy (anonymous):

u have to plug 3 in for x and -2 in for y

mathslover (mathslover):

\[\large{\color{blue}{\textbf{Put the values given in the expression i.e. x = 3 and y = -2 in }\\\textbf{x-3y}}}\] x-3y = 3-3(-2) = 3 +3(2) = 3+6=9

Parth (parthkohli):

\[\rm f(x,y) = x - 3y \\ f(3,-2) = 3 - 2(-2)\]